OpenStudy (anonymous):

The volume of a sphere is 4,500 cubic yards. What is the approximate diameter of the sphere?

OpenStudy (anonymous):

volume = (4/3) · π · r3 r^3 = 3V/4 pi r^3 = (3 x 4,500)/(4 x 3.24) r^3 = 13,500/12.56 r^3 = 1,075 r = cu rt of 1,075 or 10.24 D = 2 x 10.24 or 20.48